# Mean-field and fluctuations for hub dynamics in heterogeneous random networks

### v1

@@ 2.1.3 Proof of reduction theorem on a star
$\xi_{m}(\theta^{t}\omega,x^{t})=\zeta_{d}(t,\omega,x\omega_{{s}},x_{{s}})+\alpha\left[\frac{1}{L}\sum_{j=2}^{N}h(\theta^{t}\omega,z^{t},x_{{s},j}^{t})-\int_{\mathbb{T}}h(\theta^{t}\omega,z^{t},y)\mathrm{d}m(y)\right].$
@@
Under the hypotheses of Theorem 2.3, by choosing $D(\varepsilon,\omega)=\max\{D_{1}(\varepsilon,\omega),D_{2}(\varepsilon,\omega)\}\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}$, we estimate $\mathrm{(HFM)}$
@@
Under the hypotheses of Theorem 2.3, by choosing $D(\varepsilon,\omega)=\max\{D_{1}(\varepsilon,\omega),D_{2}(\varepsilon,\omega)\}\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}$, we control $\mathrm{(LFM)}$
@@ 2.2 Reduction on a locally star-like network
This way, the hub behavior in the $N$-dimensional network dynamics becomes reduced to (approximated by) the one-dimensional RDS $\varphi_{i,\alpha,m}:\mathbb{N}\times\Omega\times\mathbb{T}\to\mathbb{T}$ recursively defined by
$\varphi_{i,\alpha,m}(t+1,\omega,z)=\varphi_{i,\alpha,m}(1,\theta^{t}\omega,\varphi_{i,\alpha,m}(t,\omega,z))+\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\int_{\mathbb{T}}h(\theta^{t}\omega,\varphi_{i,\alpha,m}(t,\omega,z),y)\mathrm{d}m(y).$
In this notation, the hub evolution becomes
$z_{i}^{t+1}=\varphi_{i,\alpha,m}(1,\theta^{t}\omega,z_{i}^{t})+\xi_{i,m}(\theta^{t}\omega,x^{t}),~{}~{}~{}~{}i\in\mathcal{H}_{\Delta}.$
@@
$d_{\mathbb{T}}(x_{j}^{t},\varphi_{j}(t,\omega_{{s}},x_{{s},j}))\leq\varepsilon_{s}\left(\alpha\Delta_{0}^{-1}\delta\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right)~{}~{}~{}~{}\forall t\in\mathbb{N},\forall j\in\mathcal{L}_{\delta};$ (Shadowing)
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi_{j}(t,\omega_{{s}},x_{{s},j})}\xrightarrow[T\to+\infty]{*}m,~{}~{}~{}~{}\forall j\in\mathcal{L}_{\delta},$ (Typicality)
where $d_{\mathbb{T}}$ denotes the distance on the circle, the shadowing precision $\varepsilon_{s}(\delta_{s})$ is a $\mathbb{R}_{+}$-valued function converging to 0 as $\delta_{s}$ tends to 0. Then, for any error tolerance
@@
each hub $i\in\mathcal{H}_{\Delta}$ admits $(\varepsilon+\alpha(1-\nu)\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}})$-reduction to $\varphi_{i,\alpha,m}$ on initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ with exceptional asymptotic frequency at most $\rho$ with
@@
Remark 2.17.
Similar to the Reduction Theorem 2.3 on a star, we may obtain (ii) Small fluctuation in long time windows and (iii) Gaussian fluctuations as in Theorem C, provided that $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ admits shadowing intial data $(\omega_{{s}},x_{{s}})$ satisfying (Shadowing) and (Typicality) and that there is an “absolutely continous shadowing map” $(\omega,x)\mapsto(\omega_{{s}},x_{{s}})$, see Appendix A for more details.
@@ 2.2.1 Proof of reduction theorem on a locally star-like network
Proof. By Lipschitz continuity of $h(\theta^{t}\omega,z^{t},\cdot)$ and by (Shadowing), we estimate
$|\zeta_{i,d}(t,\omega,x,\omega_{{s}},x_{{s}})|\leq$ $\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}|h(\theta^{t}\omega,z_{i}^{t},\cdot)|_{\mathrm{Lip}}d_{\mathbb{T}}(x_{j}^{t},x_{{s},j}^{t})$
$\leq$ $\alpha\sup_{t\in\mathbb{N}}|h(\theta^{t}\omega,\cdot,\cdot)|_{\mathrm{Lip}}\varepsilon_{s}\left(\alpha\Delta_{0}^{-1}\delta\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right),~{}~{}~{}~{}\forall i\in\mathcal{H}_{\Delta}.$
The estimate follows by assumption of Theorem 2.16
$\varepsilon\geq 4\alpha\sup_{t\in\mathbb{N}}|h(\theta^{t}\omega,\cdot,\cdot)|_{\mathrm{Lip}}\varepsilon_{s}\left(\alpha\Delta_{0}^{-1}\delta\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right).$
@@
$D(\varepsilon,\omega)=\max\{D_{1}(\varepsilon,\omega),D_{2}(\varepsilon,\omega)\}\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}$
@@
Proof. By decay of Fourier coefficients Corollary 2.6, we estimate
$|\zeta_{i,h}(t,\omega,x,\omega_{{s}},x_{{s}})|\leq$ $\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\sum_{n_{1}\neq 0}\sum_{|n_{2}|>D}\frac{\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}}{(2\pi)^{4}n_{1}^{2}n_{2}^{2}}\frac{1}{\nu_{i}k_{i}}\nu_{i}k_{i}2\|\phi_{n_{2}}\|_{\infty}$
$+\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\sum_{|n_{2}|>D}\frac{\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}}{(2\pi)^{4}n_{2}^{4}}\frac{1}{\nu_{i}k_{i}}\nu_{i}k_{i}2\|\phi_{n_{2}}\|_{\infty}$
$\leq$ $\frac{2\alpha\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}}{(2\pi)^{4}}\sum_{n_{1}\neq 0}\frac{1}{n_{1}^{2}}\sum_{|n_{2}|>D}\frac{1}{n_{2}^{2}}+\frac{2\alpha\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}}{(2\pi)^{4}}\sum_{|n_{2}|>D}1/n_{2}^{4}.$
The estimates follow by choosing $D(\varepsilon,\omega)=\max\{D_{1}(\varepsilon,\omega),D_{2}(\varepsilon,\omega)\}$ as in Lemma 2.11. ∎
@@
$D(\varepsilon,\omega)=\max\{D_{1}(\varepsilon,\omega),D_{2}(\varepsilon,\omega)\}\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}},$
@@ 3 Examples
In this section we provide a general class of examples of node dynamics, coupling functions and network structures satsifying the dimensional reduction Theorems 2.3 and 2.16.
@@ 3.1.1 Unique stationary measure
Define the coding map by pull-back limit
@@ 3.1.2 Typical random orbits
Consider the iid random iteration $\varphi$ of uniform contractions given in (14). There is a uniform set $\Omega_{*}\subseteq\Omega$ of full $\mathbb{P}$-measure such that for every initial condition $x\in M$ and every noise $\omega\in\Omega_{*}$, the asymptotic behavior of the random orbit $\varphi(t,\omega,x)$ is described by the stationary measure
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi(t,\omega,x)}\xrightarrow[T\to+\infty]{*}\pi_{*}\mathbb{P},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi(t,\omega,x_{0})}=\frac{1}{T}\sum_{t=0}^{T-1}\delta_{Z_{t}(\omega)}\xrightarrow[T\to+\infty]{*}\pi_{*}\mathbb{P},~{}~{}~{}~{}\forall\omega\in\Omega_{x_{0}}.$
@@ 3.1.3 Shadowing and Typicality on the star
the hub behavior admits $\varepsilon$-reduction to $\varphi_{\alpha,m}$ defined in (12), for almost every noise realization $\omega$ and any initial condition $x\in\mathbb{T}^{N}$ with exponentially small exceptional asymptotic frequency at most $\rho$ with
@@
By taking $\omega\in\Omega_{*}$ as in Theorem 3.3 and any $x\in\mathbb{T}^{N}$, it suffices to verify the (Shadowing) for the low degree trajectory $\{x_{j}^{t}=\Phi_{\alpha}(t,\omega,x)_{j}:j=2,\cdots,N,t\in\mathbb{N}\}$ and its shadowing random orbit $\varphi_{j}(t,\omega,x_{j})$.
@@ 3.2 A locally star-like random power-law graph model
Proof of Theorem 3.7. I. First we compute the number $M$ of hubs by solving the equation $w_{M}=\lambda_{\mathrm{hub}}\cdot m$, which reads
@@
III. We count the number of low degree nodes among neighbors of a hub. For each hub $i=1,\cdots,M$, we consider the random variable
@@ 3.2.1 Shadowing and Typicality on a power-law network
Consider $G$-network dynamics (7) satisfying (R1–3), where $G$ is a $(\Delta,\delta,\nu)$-locally star-like network on $N\gg 1$ nodes with largest degree $\Delta_{0}$ from the random power-law graph model as in Theorem 3.7, and node dynamics given by iid random iteration of contractions $\varphi_{i}$ with uniform contraction rate $\lambda\in(0,1)$ and unique stationary measure $m_{0}$ as in Theorem 3.3. Then, for any error tolerance
@@
the $(\varepsilon+\alpha(1-\nu)\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}})$-reduction holds for all hubs, almost every noise realization $\omega$ and any initial condition $x\in\mathbb{T}^{N}$ with exponentially small exceptional asymptotic frequency at most $\rho$ with
@@
By taking $\omega\in\Omega_{*}$ and any $x\in\mathbb{T}^{N}$, it suffices to verify the (Shadowing) for the low degree trajectory $\{x_{j}^{t}=\Phi_{\alpha}(t,\omega,x)_{j}:j\in\mathcal{L}_{\delta},t\in\mathbb{N}\}$ and its shadowing random orbit $\varphi_{j}(t,\omega,x_{j})$.
@@
We have thus verified almost surely and for each initial condition $x\in\mathbb{T}^{N}$ (Shadowing) with shadowing precision $\varepsilon_{s}=\frac{1}{1-\lambda}\alpha\Delta_{0}^{-1}\delta\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}$. The corollary follows from Theorem 2.16. ∎
Appendix A Proofs of Theorems A, C, and more general settings
In addition to the hypotheses of Theorems 2.3 and 2.16, if $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ admits shadowing intial data $(\omega_{{s}},x_{{s}})$ satisfying (Shadowing) and (Typicality), then the $(m,\varepsilon)$-reduction holds $\mathbb{P}\otimes m^{\otimes N}$-almost surely. Moreover, if shadowing can be realized in an absolutely continuous way, then one can transfer the measure-theoretic results (ii) Small fluctuation in long time windows and (iii) Gaussian fluctuations from decoupled low degree node dynamics to the coupled network system.
@@ 3. $\mathrm{(iii)}$
Item (i) Almost sure reduction in asymptotic time of Theorems A.2, A and C can be proven using Corollaries 3.5 and 3.10. The particular constants are simpler and sharper in the Theorems than in the Corollaries because the coupling function in the introduction example has only one Fourier mode and hence requires no truncation; these detailed calculations are presented in full in [Bia24].

### v2

@@ 2.1.3 Proof of reduction theorem on a star
$\xi_{m}(\theta^{t}\omega,x^{t})=\zeta_{d}(t,\omega,x,\omega_{{s}},x_{{s}})+\alpha\left[\frac{1}{L}\sum_{j=2}^{N}h(\theta^{t}\omega,z^{t},x_{{s},j}^{t})-\int_{\mathbb{T}}h(\theta^{t}\omega,z^{t},y)\mathrm{d}m(y)\right].$
@@
Under the hypotheses of Theorem 2.4, by choosing $D(\varepsilon,\omega)\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}$, we estimate $\mathrm{(HFM)}$
@@
Under the hypotheses of Theorem 2.4, by choosing $D(\varepsilon,\omega)\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}$, we control $\mathrm{(LFM)}$
@@ 2.2 Reduction on a locally star-like network
$\sup_{t\in\mathbb{N}}\max_{j\in\mathcal{L}_{\delta}}d_{\mathbb{T}}(x_{j}^{t},\varphi_{j}(t,\omega_{{s}},x_{{s},j}))\leq\varepsilon_{s}\left(\alpha\Delta_{0}^{-1}\delta\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right);$ (Shadowing)
$\frac{1}{T}\sum_{t=0}^{T-1}\bigotimes_{j\in\mathcal{L}_{\delta}}\delta_{\varphi_{j}(t,\omega_{{s}},x_{{s},j})}\xrightarrow[T\to+\infty]{\text{weak}^{*}}m^{\otimes L},$ (Typicality)
where $d_{\mathbb{T}}$ denotes the distance on the circle, the shadowing precision $\delta_{s}\mapsto\varepsilon_{s}(\delta_{s})$ is a $\mathbb{R}_{+}$-valued function converging to 0 as $\delta_{s}$ tends to 0. Then, for any error tolerance
@@
each hub $i\in\mathcal{H}_{\Delta}$ admits $(\varepsilon+\alpha(1-\nu)\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}})$-reduction to $\varphi_{\alpha_{i},m}$ in Eq. (12) with $\alpha_{i}=\alpha\frac{\nu_{i}\kappa_{i}}{\Delta_{0}}$ on initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ with exceptional asymptotic frequency at most $\rho$ with
@@ 2.2.1 Proof of reduction theorem on a locally star-like network
Proof. The estimate follows by assumption of Theorem 2.18 similar to the proof of Lemma 2.12. ∎
@@
$D(\varepsilon,\omega)\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}$
@@
Proof. The estimates are similar to Lemma 2.13 with $\alpha$ replaced by $\alpha_{i}=\alpha\frac{\nu_{i}\kappa_{i}}{\Delta_{0}}$.∎
@@
$D(\varepsilon,\omega)\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}},$
@@ 3 Examples
In this section we provide a general class of examples of node dynamics, coupling functions and network structures satsifying the dimensional reduction Theorems 2.4 and 2.18.
@@ 3.1.1 Unique stationary measure
Define the coding map by pre-composition
@@ 3.1.2 Typical random orbits
Consider the iid random iteration $\varphi$ of uniform contractions given in (14) satisfying (C1-4). There is a uniform set $\Omega_{*}\subseteq\Omega$ of full $\mathbb{P}$-measure such that for every initial condition $x\in M$ and every noise $\omega\in\Omega_{*}$, the asymptotic behavior of the random orbit $\varphi(t,\omega,x)$ is described by the stationary measure
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi(t,\omega,x)}\xrightarrow[T\to+\infty]{\text{weak}^{*}}\pi_{*}\mathbb{P},$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi(t,\omega,x_{0})}=\frac{1}{T}\sum_{t=0}^{T-1}\delta_{Z_{t}(\omega)}\xrightarrow[T\to+\infty]{\text{weak}^{*}}\pi_{*}\mathbb{P},~{}~{}~{}~{}\forall\omega\in\Omega_{x_{0}}.$
@@ 3.1.3 Shadowing and Typicality on the star
the hub behavior admits $\varepsilon$-reduction to $\varphi_{\alpha,m_{0}}$ defined in (12), for almost every noise realization $\omega$ and any initial condition $x\in\mathbb{T}^{N}$ with exponentially small exceptional asymptotic frequency at most $\rho$ with
@@
By (R1) and (C4), the $\varphi_{i}$ are independent random uniform contractions on $\mathbb{T}$ with rate $\lambda$. It follows that the uncoupled system $\Phi_{\alpha=0}$ is a random uniform contraction satisfying (C4) on $\mathbb{T}^{N}$, equipped with the distance $d_{\mathbb{T}^{N}}(x,y)=\max_{i=1,\cdots,N}d_{\mathbb{T}}(x_{i},y_{i})$, with the same rate $\lambda$. By Theorem 3.3 applied to $\Phi_{\alpha=0}$ on $(\mathbb{T}^{N},d_{\mathbb{T}^{N}})$, the random orbit $\Phi_{\alpha=0}(t,\omega,x)$ starting from any $x\in\mathbb{T}^{N}$ and $\omega\in\Omega_{*}$ asymptotically distributes as the unique stationary measure $m_{0}^{\otimes N}$.
To verify the (Typicality) assumption in Theorem 2.4, simply take $x_{s}=x\in\mathbb{T}^{N}$ and $\omega_{s}=\omega\in\Omega_{*}$. For the (Shadowing) assumption, we compute
@@ 3.2 A locally star-like random power-law graph model
Proof of Theorem 3.7. Our strategy is to count the number of nodes that are expected to be hubs in Step I and low degree nodes in Step II, prove the expected picture satisfies the locally star-like property in Step III, and finally bring the estimates to the actual degrees by concentration.
I. First we compute the number $M$ of expected hubs by solving the equation $w_{M}=\lambda_{\mathrm{hub}}\cdot m$, which reads
@@
III. To count the number of expected low degree nodes among neighbors of an expected hub $i=1,\cdots,M$, we consider the random variable
@@ 3.2.1 Shadowing and Typicality on a power-law network
Consider $G$-network dynamics (7) satisfying (R1–3), where $G$ is a $(\Delta,\delta,\nu)$-locally star-like network on $N\gg 1$ nodes with largest degree $\Delta_{0}$ from the random power-law graph model as in Theorem 3.7, and node dynamics given by iid random iteration of contractions $\varphi_{i}$ with uniform contraction rate $\lambda\in(0,1)$ and unique stationary measure $m_{0}$ as in Corollary 3.5. Then, for any error tolerance
@@
each hub $i\in\mathcal{H}_{\Delta}$ admits $(\varepsilon+\alpha(1-\nu)\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}})$-reduction to $\varphi_{\alpha_{i},m_{0}}$ in Eq. (12) with $\alpha_{i}=\alpha\frac{\nu_{i}\kappa_{i}}{\Delta_{0}}$, for almost every noise realization $\omega$ and any initial condition $x\in\mathbb{T}^{N}$, with exceptional asymptotic frequency at most $\rho$ with
@@
As in the proof of Corollary 3.5, we take $x_{s}=x\in\mathbb{T}^{N}$ and $\omega_{s}=\omega\in\Omega_{*}$, verifying the (Typicality) assumption in Theorem 2.18. For the (Shadowing) assumption, we compute
@@
We have thus verified almost surely and for each initial condition $x\in\mathbb{T}^{N}$ the (Shadowing) assumption with shadowing precision $\varepsilon_{s}=\frac{1}{1-\lambda}\alpha\Delta_{0}^{-1}\delta\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}$. The corollary follows from Theorem 2.18. ∎
Appendix A Proofs of Theorems A, C and more general settings
In addition to the hypotheses of Theorems 2.4 and 2.18, if $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ admits shadowing intial data $(\omega_{{s}},x_{{s}})$ satisfying (Shadowing) and (Typicality), then the $(m,\varepsilon)$-reduction holds $\mathbb{P}\otimes m^{\otimes N}$-almost surely. Moreover, if shadowing can be realized in an absolutely continuous way, then one can transfer the measure-theoretic results (ii) Small fluctuation in long time windows and (iii) Gaussian fluctuations from decoupled low degree node dynamics to the coupled network system.
@@ 3. $\mathrm{(iii)}$
Item (i) Almost sure reduction in asymptotic time of Theorem A.2 can be proven using Corollaries 3.5 and 3.10. The particular constants are simpler and sharper in the Theorem than in the Corollaries because the coupling function in the introduction example has only one Fourier mode and hence requires no truncation; these detailed calculations are presented in full in [Bia24].
@@
A.1 Proof of Theorem A
Since the coupling function (3) has only one Fourier mode, the Fourier truncation in Lemma 2.13 is unnecesary and the bad set estimates in Lemma 2.14 reduce to one bad set only.
Item (i) follows by a simplified version of Theorem 2.4 together with the uniform typicality Theorem 3.3. The shadowing precision $\varepsilon_{s}=\frac{17\alpha}{3L}$ can be calculated as in Corollary 3.5. We choose $\varepsilon_{b}=\frac{\varepsilon}{2\alpha}$ so that the shadowing and bad set arguments each contribute $\varepsilon/2$ to the reduction estimate.
Item (ii) follows from a similar argument to Theorem A.2 Item (ii) with $\varepsilon_{b}=L^{-\kappa}\sqrt{6}$ and $\varepsilon_{s}=\frac{17\alpha}{3L}$.
Item (iii) follows from Theorem A.2 Item (iii) with $\phi(x)=\sin 2\pi x$, $\psi(\omega)=\omega/3.6$.
A.2 Proof of Theorem C
Item (i) follows by a simplified version of Theorem 2.18 together with the uniform typicality Theorem 3.3. The non ldn contribution from Lemma 2.20 is $|\zeta_{i,c}|\leq\frac{17}{6}\alpha(1-\nu)\leq\varepsilon/3$, by our choice of $\varepsilon$. The shadowing precision $\varepsilon_{s}=\frac{17\alpha\delta}{3\Delta}$ can be calculated as in Corollary 3.10. We choose $\varepsilon_{b}=\frac{\varepsilon}{3\alpha}$ so that the shadowing and bad set arguments each contribute $\varepsilon/3$ to the reduction estimate.
Item (ii) follows from a similar argument to Theorem A.2 Item (ii) with $\varepsilon_{b}=\Delta^{-\kappa}\sqrt{6}$ and $\varepsilon_{s}=\frac{17\alpha\delta}{3\Delta}$.
Item (iii) follows from a similar argument to Theorem A.2 Item (iii) with $\phi(x)=\sin 2\pi x$, $\psi(\omega)=\omega/3.6$. The correction constants $s_{1},s_{2}$ need to account for non ldn contribution.
$s_{1}:=\alpha_{i}\frac{1}{\nu_{i}k_{i}}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}(\sin 2\pi x_{i}^{t}-\sin 2\pi y_{j}^{t})+\alpha\frac{1}{\Delta_{0}}\sum_{j\in\mathcal{N}\setminus\mathcal{L}_{\delta}}h_{\omega_{i}^{t}}(z_{i}^{t},x_{j}^{t}).$
Using $|h|\leq 2+\frac{3}{3.6}=\frac{17}{6}$ and $\varepsilon_{s}=\frac{17}{3}\alpha\frac{\delta}{\Delta}$, we have
$|s_{1}|\leq\alpha_{i}2\pi\varepsilon_{s}+\alpha\Delta_{0}^{-1}(1-\nu_{i})k_{i}\frac{17}{6}=\alpha_{i}2\pi\frac{17}{3}\alpha\frac{\delta}{\Delta}+\alpha\Delta_{0}^{-1}(1-\nu_{i})k_{i}\frac{17}{6},$
which simplifies to
$|s_{1}|=O(N^{-\frac{1-\lambda_{\mathrm{ldn}}}{\beta-1}}+N^{-\lambda_{\mathrm{ldn}}\frac{\beta-1}{\beta-2}}w^{\beta-2}).$
Lastly, the Berry-Esseen correction, similar to that in Theorem A.2 Item (iii), reads
$s_{2}:=\frac{C\int_{0}^{1}|\sin 2\pi x|^{3}\mathrm{d}x}{2^{-3/2}\sqrt{\nu_{i}k_{i}}}=O(N^{-\frac{1}{2\beta-2}}),~{}~{}~{}~{}\forall i\in\mathcal{H}_{\Delta}.$